A man travel 128km at average speed of x km/h he find that he could have save 9/2 hours if he had increased his speed by 2 km/h.find x
Answer by mananth(16946)

speed = x kmph
time = distance / speed
= 128/x
new speed = x+2
time = 128/ x+2
128/x - 128/ x+2 = 9/2
LCD = x(x+2)
128(x+2)-128x / x(x+2)= 9/2
128x+256-128x/x(x+2)= 9/2
256 = 9x(x+2)/2
2*256= 9x^2+18x
512=9x^2+18x
9x^2+18x-512=0
find the roots x1,x2
a=9, b=18, c=-512
x1= (-18+sqrt(18^2+4*9*512))/18
x = 6.6
x2=(-18-sqrt(18^2+4*9*512))/18
x2=-8.6 ignore this negative value.